OpenStudy (anonymous):

2.5 qt/min = ____ gal/h

OpenStudy (mrnood):

there are 4 quarts in a gallon (at least in the UK there are...) There are 60 mins in 1 hour Please work it out from there....

OpenStudy (mathmale):

\[\rightarrow \frac{ 2.5 qt }{ \min }*\frac{ 60 \min }{ 1 hr }*\frac{ 1 gal }{ 4 qt }\]

OpenStudy (mathmale):

Please try simplifying that. What's your final result?

OpenStudy (mrnood):

@mathmale Does it help to give anyone a complete solution? If you don't have to think about anything you never learn anything.

OpenStudy (mathmale):

I owe you an apology. I didn't look carefully to see whether our student had listed those equivalents himself/herself. Since you listed them, I should have started by asking the student to come up with the appropriate conversion units.

OpenStudy (mrnood):

Sorry - I see from your profile that you are way more into the academic world than I. I just get frustrated at the number of Qs here which demand an answer to a simple equation. Sorry if I was a little abrupt!

OpenStudy (mathmale):

That's OK; think nothing of it. I share your discomfort over all those requests for answers with no work involved on the part of the "learner." All the best to you.